OpenStudy (anonymous):

If 5^x = b, evaluate (25)^x * (125)^3x

OpenStudy (anonymous):

Here \[5^{x}=b\] \[25^{x}\times125^{3x}\] \[(5^{2})^{x}(5^{3})^{3x})\] \[5^{2x}\times5^{9x}\] \[5^{2x+9x}\] \[5^{11x}\] \[(5^{x})^{11}\] \[b ^{11}\]

OpenStudy (anonymous):

Is it Ok

OpenStudy (anonymous):

If 5^x = b, evaluate (25)^x * (125)^3x (25)^x * (125)^3x =(5*5)^x (5*5*5*5)^3x =(5^2)^x (5^3)^3x =(5)^2x (5)^9x =5^11x =(5^x)^11 =b^11